Grade 12th passMechanics

a bullet travelling with velocity 100m/s is able to penetrate planck 4cm thick.T he velocity of the ullet to penetrate into planck of 25 cm thickness is

rojasree
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
The energy required to pierce a 4 cm plank is (1/2) * m * v^2 = (1/2) * m * (100 m/s)^2 = (5000 m^2/s^2) 8 m. The amount of energy required to pierce a 25 cm plank will be 25/4 as much, or 11250 * m. 11,250 * m = (1/2) * m * v^2, or v = sqareroot(22,500) = 150 m/s.